OpenStudy (anonymous):

When the sum of a number plus 3 is squared, it is 9 more than the sum of the number plus 2 when squared. @Rex831 @matty546

OpenStudy (anonymous):

x=2

OpenStudy (anonymous):

(x + 3)^2 - 9 = (x + 2)^2 x^2 + 6x + 9 - 9 = x^2 + 4x + 4 2x = 4 x = 2

OpenStudy (anonymous):

All good now, @shortie212 ?

OpenStudy (anonymous):

yea i posted another one @tcarroll010